OpenStudy (anonymous):

Which of the following are solutions to the equation below? x2 + 16 = -10x

OpenStudy (blurbendy):

x^2 + 10x + 16 = 0 (x + 8)(x + 2) = 0 x = -8 x = -2

OpenStudy (anonymous):

Thank you